# Designing Decision Support Systems Using Counterfactual Prediction Sets

### v2

@@ 6 Discussion and Limitations
In this section, we discuss several assumptions and limitations of our work, pointing out interesting avenues for future research.
Data. We have assumed that the data samples and the expert predictions are drawn i.i.d. from a fixed distribution and the calibration set contains samples with noiseless ground truth labels. In future work, it would be very interesting to lift these assumptions. Regarding distribution shift, a good starting point may be the rapidly expanding literature on conformal prediction under distribution shift [18, 47, 35]; regarding label noise, the recent work by [16], which has found that conformal prediction is robust to label noise; and, regarding label ambiguity, the very recent work by [45].
@@
Decision making task. We have focused on designing decision support systems for multiclass classification tasks. It would be very interesting to extend our approach and our notion of counterfactual monotonocity to other classification tasks. To this end, a good starting point may be the framework of risk controlling prediction sets (RCPS) by [5], which generalizes conformal prediction and does allow for a variety of classification tasks, including multilabel classification. Further, it would also be very interesting to investigate to what extend our ideas are useful in other types of decision tasks (e.g., reinforcement learning) and decision support systems (e.g., LLMs).
@@ 7 Conclusions
We have looked at the development of decision support systems based on prediction sets for multiclass classification tasks from the perspective of online learning and counterfactual inference. This perspective has allowed us to design a methodology that does not require, nor assumes, a stylized human expert model, and has modest computational and data requirements. In doing so, we have identified two natural monotonicity assumptions which may be of independent interest. Further, we have conducted a human subject study, which demonstrates that our methodology is superior to the state of the art and adaptively limiting experts’ level of agency may lead to greater performance.
Broader Impact
Our work has focused on maximizing the average accuracy achieved by a human expert using a decision support system based on conformal prediction. However, as discussed in Section 6, whenever the expert’s predictions are consequential to individuals—and as a result the society—, it would be important to extend our methodology to account for fairness considerations. Within Section 6, we have also discussed other assumptions and limitations of our work, which may influence its impact in practice.
@@ Appendix A Proof of Theorem 1
where $\bar{\mathcal{E}}$ is the complement of $\mathcal{E}$. Then, following Slivkins, [42], we first assume =$\mathcal{E}=\{|\hat{\mu}_{t}(\alpha)-\mathbb{E}[\mathbb{I}\{\hat{Y}_{\mathcal{C}_{\alpha}}=Y\}]|\leq\epsilon_{t}(\alpha)|,\forall\alpha\in\mathcal{A},\forall t\leq T\}$ holds and then show that the probability that $\bar{\mathcal{E}}$ holds is negligible.
@@
From the above, it immediate follows that anytime we pull $\alpha$, we suffer average regret $\Delta(\alpha)$. Consequently, for any time $t\leq T$, the total average regret due to pulling arm $\alpha$, which we denote as $\mathbb{E}[R(\alpha\,;\,t)\,|\,\mathcal{E}]$, is given by $\mathbb{E}[R(\alpha\,;\,t)\,|\,\mathcal{E}]=n_{t}(\alpha)\cdot\Delta(\alpha)$, where $n_{t}(\alpha)$ denotes the number of times that arm $\alpha$ has been pulled until $t$. Thus, for any time $t\leq T$, the total average regret conditioned on $\mathcal{E}$ is given by:
@@ Appendix B Additional Details about the Human Subject Study Setup
Dataset. The ImageNet16H dataset [43] was created using $1{,}200$ unique images labeled into $207$ different fine-grained categories from the ImageNet Large Scale Visual Recognition Challenge (ILSRVR) 2012 dataset [37]. More specifically, in the ImageNet16H dataset, each of the above images was used to generate four noisy images, each with a different amount of phase noise distorsion $\omega\in\{80,95,110,125\}$, and each of the above fine-grained categories was mapped into one out of $16$ coarse-grained categories (i.e., chair, oven, knife, bottle, keyboard, clock, boat, bicycle, airplane, truck, car, elephant, bear, dog, cat, and bird), which serve as ground truth labels [^15] Note that, by mapping the fine-grained categories used in the ILSRVR 2012 dataset into coarse-grained categories, one essentially gets rid of any potential label disagreement among annotators in the ILSRVR 2012 dataset.. The amount of phase noise distortion controls the difficulty of the classification task; the higher the noise, the more difficult the classification task. In our experiments, we used all $1{,}200$ noisy images with $\omega=110$ because, under such noise value, humans sometimes, but not always, succeed at solving the prediction task (i.e., the empirical success probability achieved by the human experts on their own was $0.760$).
@@ Appendix D Sensitivity Analysis to Violations of the Counterfactual Monotonicity Assumption
Experimental setup. Since counterfactual monotonicity lies within level three in the “ladder of causation” [34], we cannot directly increase (nor estimate!) the amount of counterfactual monotonicity violations in the experts’ predictions gathered in our human subject study. However, we can increase the amount of interventional monotonicity violations (i.e., how frequently Eq. 3 is violated) and, since interventional monotonicity is a necessary condition for counterfactual monotonicity, we argue that we are indirectly increasing the amount of counterfactual monotonicity violations. To this end, we randomly select a fraction $p_{v}$ of the images used in our human subject study and, for each of these images, we randomly permute the values $\mathbb{I}\{\hat{y}_{\mathcal{C}_{\alpha}}=y\}$ across pairs of experts’ predictions $\hat{y}_{\mathcal{C}_{\alpha}}$ and prediction sets $\mathcal{C}_{\alpha}(x)$ such that $y\in\mathcal{C}_{\alpha}(x)$. Here, the larger the fraction $p_{v}$ of images we permute their labels, the larger the amount of interventional (counterfactual) monotonicity violations, as shown in Figure 7. In the Figure, we stratify the images into groups of similar difficulty, following the procedure described in Appendix B, to demonstrate that the effect of the above random permutations is more apparent in the images of higher difficulty. [^16] We do not show the empirical success probability for images of low difficulty because it is originally almost always $1$ and thus it is not affected by the random permutations.

### v3

@@ 6 Discussion and Limitations
In this section, we discuss several assumptions and limitations of our work, pointing out avenues for future research.
Data. We have assumed that the data samples and the expert predictions are drawn i.i.d. from a fixed distribution and the calibration set contains samples with noiseless ground truth labels. In future work, it would be very interesting to lift these assumptions. Regarding distribution shift, a good starting point may be the rapidly expanding literature on conformal prediction under distribution shift (Tibshirani et al.,, 2020; Podkopaev and Ramdas,, 2021; Gibbs and Candes,, 2021); regarding label noise, the recent work by Einbinder et al., (2022), which has found that conformal prediction is robust to label noise; and, regarding label ambiguity, the very recent work by Stutz et al., (2023).
@@
Decision making task. We have focused on designing decision support systems for multiclass classification tasks. It would be very interesting to extend our approach and our notion of counterfactual monotonocity to other classification tasks. To this end, a good starting point may be the framework of risk controlling prediction sets (RCPS) by Bates et al., (2021), which generalizes conformal prediction and does allow for a variety of classification tasks, including multilabel classification. Further, it would also be very interesting to investigate to what extend our ideas are useful in other types of decision tasks (e.g., reinforcement learning) and decision support systems (e.g., LLMs).
@@ 7 Conclusions
We have looked at the development of decision support systems based on prediction sets for multiclass classification tasks from the perspective of online learning and counterfactual inference. This perspective has allowed us to design a methodology that does not require, nor assumes, a stylized human expert model, and has modest computational and data requirements. In doing so, we have also identified two natural monotonicity assumptions, i.e., interventional monotonicity and counterfactual monotonicity, which may be of independent interest. Further, we have conducted a large-scale human subject study that shows that our methodology is superior to the state of the art and, for decision support systems based on prediction sets, adaptively limiting experts’ level of agency leads to greater performance.
@@ Acknowledgements
Impact Statement
Our work has focused on maximizing the average accuracy achieved by a human expert using a decision support system based on conformal prediction. However, whenever the expert’s predictions are consequential to individuals, it would be important to extend our methodology to account for fairness considerations. In Section 6, we have also discussed other assumptions and limitations of our work, which may influence its impact in practice.
@@ Appendix A Proof of Theorem 1
where $\bar{\mathcal{E}}$ is the complement of $\mathcal{E}$. Then, following Slivkins, (2019), we first assume =$\mathcal{E}=\{|\hat{\mu}_{t}(\alpha)-\mathbb{E}[\mathbb{I}\{\hat{Y}_{\mathcal{C}_{\alpha}}=Y\}]|\leq\epsilon_{t}(\alpha)|,\forall\alpha\in\mathcal{A},\forall t\leq T\}$ holds and then show that the probability that $\bar{\mathcal{E}}$ holds is negligible.
@@
From the above, it immediately follows that anytime we pull $\alpha$, we suffer average regret $\Delta(\alpha)$. Consequently, for any time $t\leq T$, the total average regret due to pulling arm $\alpha$, which we denote as $\mathbb{E}[R(\alpha\,;\,t)\,|\,\mathcal{E}]$, is given by $\mathbb{E}[R(\alpha\,;\,t)\,|\,\mathcal{E}]=n_{t}(\alpha)\cdot\Delta(\alpha)$, where $n_{t}(\alpha)$ denotes the number of times that arm $\alpha$ has been pulled until $t$. Thus, for any time $t\leq T$, the total average regret conditioned on $\mathcal{E}$ is given by:
@@ Appendix B Additional Details about the Human Subject Study Setup
Dataset. The ImageNet16H dataset (Steyvers et al.,, 2022) was created using $1{,}200$ unique images labeled into $207$ different fine-grained categories from the ImageNet Large Scale Visual Recognition Challenge (ILSRVR) 2012 dataset (Russakovsky et al.,, 2015). More specifically, in the ImageNet16H dataset, each of the above images was used to generate four noisy images, each with a different amount of phase noise distorsion $\omega\in\{80,95,110,125\}$, and each of the above fine-grained categories was mapped into one out of $16$ coarse-grained categories (i.e., chair, oven, knife, bottle, keyboard, clock, boat, bicycle, airplane, truck, car, elephant, bear, dog, cat, and bird), which serve as ground truth labels [^15] Note that, by mapping the fine-grained categories used in the ILSRVR 2012 dataset into coarse-grained categories, one essentially gets rid of any potential label disagreement among annotators in the ILSRVR 2012 dataset.. The amount of phase noise distortion controls the difficulty of the classification task; the higher the noise, the more difficult the classification task. In our experiments, we used all $1{,}200$ noisy images with $\omega=110$ because, under such noise value, humans sometimes, but not always, succeed at solving the prediction task (i.e., the empirical success probability achieved by the human experts on their own was $0.760$).
@@ Appendix D Sensitivity Analysis to Violations of the Counterfactual Monotonicity Assumption
Experimental setup. Since counterfactual monotonicity lies within level three in the “ladder of causation” (Pearl,, 2009), we cannot directly increase (nor estimate!) the amount of counterfactual monotonicity violations in the experts’ predictions gathered in our human subject study. However, we can increase the amount of interventional monotonicity violations (i.e., how frequently Eq. 3 is violated) and, since interventional monotonicity is a necessary condition for counterfactual monotonicity, we argue that we are indirectly increasing the amount of counterfactual monotonicity violations. To this end, we randomly select a fraction $p_{v}$ of the images used in our human subject study and, for each of these images, we randomly permute the values $\mathbb{I}\{\hat{y}_{\mathcal{C}_{\alpha}}=y\}$ across pairs of experts’ predictions $\hat{y}_{\mathcal{C}_{\alpha}}$ and prediction sets $\mathcal{C}_{\alpha}(x)$ such that $y\in\mathcal{C}_{\alpha}(x)$. Here, the larger the fraction $p_{v}$ of images we permute their labels, the larger the amount of interventional (counterfactual) monotonicity violations, as shown in Figure 7. In the Figure, we stratify the images into groups of similar difficulty, following the procedure described in Appendix B, to demonstrate that the effect of the above random permutations is more apparent in the images of higher difficulty. [^16] We do not show the empirical success probability for images of low difficulty because it is originally almost always $1$ and thus it is not affected by the random permutations.